Solve. Please show all of your work.
If the first and third of three consecutive odd integers are added, the result is 69 less than five times the second integer. Find all three integers.

Answer by stanbon(75887)

If the first and third of three consecutive odd integers are added, the result is 69 less than five times the second integer. Find all three integers.
-----------
1st: 2x-1
2nd: 2x+1
3rd: 2x+3
-----------
Equation:
2x-1+2x+3 = 5(2x+1)-69
--------
4x+2 = 10x-64
6x = 66
x = 11
---
1st: 2x-1 = 21
2nd: 2x+1 = 23
3rd: 2x+3 = 25
=======================
Cheers,
Stan H.